# Some frameworks for dissipative evolution in multiscale non-equilibrium thermodynamics

### v1

@@ 6.3 Contact structure geometry
We have seen that by adding dissipation to the microscopic Hamiltonian dynamics, we bring attention to overall features of the dynamics that are seen in our everyday observations of the behavior of macroscopic systems. The pure nondissipative Hamiltonian dynamics as well as the pure dissipative dynamics are two extreme idealizations of our experience. But dynamics at these two extremes is, from the mathematical point of view, understood best. In the former case the fundamental group of the dynamics is the group of canonical transformation (transformations preserving symplectic structure) and in the latter case it is the group of Legendre transformations (transformations arising in maximization of entropy in the presence of constraints).
@@
Following the Lagrange strategy, we begin the process of bringing micro and macro dynamics to a harmony on the side of macro (i.e., in classical equilibrium thermodynamics). We consider the conjugate state variables $\mathbf{x}^{*}$ as quantities playing the role of Lagrange multipliers and then, together with the entropy $s$ that is the maximized potential, we adopt them as extra state variables. The additional new step now is to realize that such an enlarged space is naturally equipped with the contact structure defined by the 1-form $\eta=ds-\mathbf{x}^{*}d\mathbf{x}$. This 1-form also has, in the context of classical equilibrium thermodynamics, an important physical significance. It can be shown [32, 34, 17, 18] that the GENERIC dynamics becomes, on the Legendre manifold (i.e., a submanifold in the enlarged space on which the 1-form $\eta=0$), dynamics preserving the contact structure.
@@ 7 Conclusion
We have reviewed several frameworks for dissipation in non-equilibrium thermodynamics, highlighting their key features and interconnections. Each framework offers unique advantages and is suited for different types of problems in thermodynamics.
@@ Acknowledgment
MP was supported by Czech Science Foundation, project 23-05736S. MP is a member of the Nečas Centre for Mathematical Modeling. We are grateful to Oğul Esen and François Gay-Balmaz for valuable discussions.

### v2

@@ 8 Contact structure geometry
We have seen that by adding dissipation to the microscopic Hamiltonian dynamics, we bring attention to overall features of the dynamics that are seen in our everyday observations of the behavior of macroscopic systems. The pure nondissipative Hamiltonian dynamics as well as the pure dissipative dynamics are two extreme idealizations of our experience. But the dynamics at these two extremes are, from the mathematical point of view, understood best. In the former case the fundamental group of the dynamics is the group of canonical transformation (transformations preserving symplectic structure) and in the latter case it is the group of Legendre transformations (transformations arising in maximization of entropy in the presence of constraints).
@@
Following the Lagrange strategy, we begin the process of bringing micro and macro dynamics to a harmony on the side of macro (i.e., in classical equilibrium thermodynamics). We consider the conjugate state variables $\mathbf{x}^{*}$ as quantities playing the role of Lagrange multipliers and then, together with the entropy $s$ that is the maximized potential, we adopt them as extra state variables. The additional new step now is to realize that such an enlarged space is naturally equipped with the contact structure defined by the 1-form $\eta=ds-\mathbf{x}^{*}d\mathbf{x}$. This 1-form also has, in the context of classical equilibrium thermodynamics, an important physical significance. It can be shown [45, 48, 82, 26, 27] that the GENERIC dynamics becomes, on the Legendre manifold (i.e., a submanifold in the enlarged space on which the 1-form $\eta=0$), dynamics preserving the contact structure.
@@ 9 Conclusion
We have reviewed several geometric frameworks for dissipation in non-equilibrium thermodynamics, highlighting their key features and interconnections. Each framework offers unique advantages and is suited for different types of problems in thermodynamics.
@@ Acknowledgment
MP was supported by Czech Science Foundation, project 23-05736S. MP is a member of the Nečas Center for Mathematical Modeling. MP used Github Copilot for grammar cleaning and for minor code suggestions. We are grateful to Oğul Esen and François Gay-Balmaz for valuable discussions on differential geometry and its applications in non-equilibrium thermodynamics. We also thank Mark Peletier and Marcello Seri for discussing the statistical background of gradient dynamics.